# SafeGen: Mitigating Unsafe Content Generation in Text-to-Image Models

### v1

@@ 6.4.1 Different Hyperparameters of Loss Weights
Diffusion Scheduler
NRR (%) $\uparrow$
CLIP
Score $\downarrow$
LPIPS
Score $\downarrow$
CLIP
Score $\uparrow$
Adversarial Prompts
(SneakyPrompt-Natural)
Benign Prompts
(COCO-25k)
PNDM 95.2% 17.85 0.789 24.60
LMSD 96.6% 18.31 0.792 24.40
EulerDiscrete 96.6% 18.52 0.794 24.42
EulerAncestorDiscrete 94.2% 17.98 0.811 24.54
DPMSovler 93.7% 18.70 0.786 24.44
DDPM 94.7% 17.97 0.810 24.40
DDIM 91.3% 18.50 0.789 24.49
Table VI: [RQ4-2] Performance of SafeGen using different diffusion schedulers during inference.
6.4.2 Different Diffusion Schedulers
Table VI shows the performance of SafeGen when using different diffusion schedulers. We assess seven prevalent schedulers [49], each with unique trade-offs between denoising speed and quality. Our intent is to quantitatively measure which scheduler works best for our use case. While all schedulers perform well in terms of NSFW removal and benign preservation under 50 denoising steps, there are slight differences between these two perspectives. For NSFW removal, the NRRs of PNDM, LMSD, and EulerDiscrete all surpass 95.0%, while PNDM has a slightly lower CLIP score indicating better breaking of the text-to-image alignment for adversarial prompts. For benign preservation, PNDM has a comparable LPIPS score with other schedulers while hits the highest CLIP score, delivering the most significant CLIP gap of 6.75 between two opposite sub-tasks. Therefore, LMSD and EulerDiscrete would be the diffusion schedulers of choice when the primary goal is to remove the NSFW content, while in this paper, we set PNDM as the default scheduler because it achieves the optimal balance between NSFW removal and benign preservation.
6.4.3 Different Denoising Diffusion Steps
Table VII demonstrates the performance of SafeGen employing different denoising diffusion steps with the default PNDM scheduler. The number of steps has been shown to be a vital hyperparameters that affects the fidelity of generated images [50]. This experiment aims to determine a proper value benefits both NSFW removal and benign preservation of SafeGen. In terms of NSFW removal, the 25- and 50-step groups offer high NRRs and low CLIP scores that denote their effective disruption of creating images that align with adversarial prompts. For benign prompts, the 50-step group also offers the highest CLIP score with comparable LPIPS score to other groups. Consequently, our choice of default denoising steps is 50, which satisfies both objectives. It is also noteworthy that less steps offering faster image creation speed to meet with practical requirements.
Table VII: [RQ4-3] Performance of SafeGen with different denoising diffusion steps during inference.
[t]
Steps
NRR (%) $\uparrow$
CLIP
Score $\downarrow$
LPIPS
Score $\downarrow$
CLIP
Score $\uparrow$
Adversarial Prompts
(SneakyPrompt-Natural)
Benign Prompts
(COCO-25k)
25 98.6% 17.12 0.794 24.16
50 95.2% 17.85 0.789 24.58
75 90.3% 18.47 0.786 24.53
100 96.6% 18.73 0.792 24.47
125 90.3% 18.62 0.792 24.46
150 90.8% 19.37 0.776 24.50
Table VII: [RQ4-3] Performance of SafeGen with different denoising diffusion steps during inference.
7 Related Work
7.1 Attacks on Text-to-Image (T2I) Models
The susceptibility of T2I models to generating NSFW content, particularly explicit nudity, has been a significant concern [7, 51, 52]. This issue has spurred investigations into various attack vectors targeting these models, such as red-teaming the SD model for unsafe image generation [53] through reverse engineering its safety filter mechanism. Moreover, adversarial prompts [54, 16, 55] were crafted to manipulate T2I models into producing NSFW images, while circumventing detection mechanisms. For instance, Ring-A-Bell [54] tailors adversarial textual inputs that are conceptually close to the target yet contain nonsense words. Gao et al. [55] introduced a word-level similarity constraint to mimic realistic human errors, e.g., typo, glyph, and phonetic mistakes. SneakyPrompt [16] employs a reinforcement learning-based search approach to craft adversarial prompts, which is effective in maintaining NSFW semantics and alluring SD models to generate unsafe images. Another vulnerability of T2I models is their reliance on large datasets, which are prone to poisoning attacks. Adversaries may release poisoned text-image data on the internet [56], which are then inadvertently collected by data trainers, leading to potential unethical outputs from T2I models.
7.2 Securing Text-to-Image (T2I) Models
The generation of pornographic content has highlighted the critical need to regulate T2I models. Current strategies focus on filtering out inappropriate textual inputs or visual outputs and mitigating textually pornographic concepts.
External mitigation measures employ plug-in safety filters [10, 11] without modifying the T2I model. The main technique is to detect unsafe content during inference or to filter out NSFW data for retraining [12]. Such filters [10, 11] protect users from exposure to inappropriate content. Similarly, SD-V2.1 [12] undergoes complete retraining on censored data using these filters. Nonetheless, existing external approaches suffer from under-generalization and susceptibility to adversarial examples. Internal mitigation methods alter the T2I model to restrain pornographic textual cues that guide image generation. Existing internal methods are mostly text-dependent. The guidance-based SLD [13] defines prohibited concepts and steers the classifier-free guidance away from unsafe regions with a new conditioned item. ESD [14, 57] fine-tunes the model, where terms like “nudity” and “sexy” are replaced with “[blank]”. Unfortunately, these methods have difficulty nullifying connotations in the text input, e.g., the names of porn stars, which are challenging to enumerate and erase.
Our Approach. Unlike prior countermeasures, SafeGen makes the first attempt to remove representations of visually sexual content from Stable Diffusion in a text-agnostic manner. This effectively cuts off the link between sexually connoted text and visually explicit content. In addition, SafeGen retains the capability for benign image generation and can seamlessly integrate with existing defense techniques. The efficacy of our design is substantiated through extensive experiments.
8 Discussion and Conclusion
@@
In this paper, we delve into the critical misuse of text-to-image (T2I) models in generating sexually explicit images. To address this risk, we introduce SafeGen, a novel framework that effectively eliminates latent representations of nudity within T2I models while preserving the models’ capability to produce high-fidelity benign content, by regulating the vision-only self-attention layers. SafeGen severs the associations between explicit visual representations and conceptually sexual prompts. As a result, it outperforms eight baselines across four datasets and achieves optimal efficacy by complementing other techniques, reaching a 99.1% NSFW removal rate (NRR).
Limitation: Our work is limited by the absence of user involvement in explicit content-related experiments due to ethical considerations. The statistics of NSFW removal rate are conducted using NudeNet [44], which might result in measurement deviations across experiment outcomes. Our manual validation partially redressed this problem by demonstrating that the results conform with the relative efficacy of different mitigation strategies.
Future Works: This work aims to shed light on model governance and promote responsible AI. We are dedicated to further contributing to the community in the following three aspects:
  • Responsible Disclosure. We have constructed a benchmark with extensive adversarial prompts. Out of ethical consideration, we provide these samples upon request to researchers for research-only purposes with institutional approvals.
  • Community Contribution. We open-source our implementation [20] and call for awareness of model compliance. We plan to promote the integration of SafeGen into widely used generative model libraries, e.g., Diffusers [49].
  • Broader Application. We envision our vision-only regulation can be extended to other generative models, including text-to-video and image-to-image models, to prevent the generation of NSFW in these applications.
@@ Appendix A Adversarial Prompts in §3
Appendix B Proof of Loss Mosaic
$z^{n}_{T}-\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}\rightarrow z^{m}_{0}$ (7)
$z^{n}_{T}=z^{n}_{0}+\sum_{t=0}^{T}{\epsilon^{n}_{t}}$ (8)
$z^{m}_{T}=z^{m}_{0}+\sum_{t=0}^{T}{\epsilon^{m}_{t}}$ (9)
$\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}\rightarrow z^{n}_{T}-z^{m}_{0}$ (10)
$\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}\rightarrow z^{n}_{T}-(z^{m}_{T}-\sum_{t=0}^{T}{\epsilon^{m}_{t}})$ (11)
$\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}-(z^{n}_{T}-z^{m}_{T}+\sum_{t=0}^{T}{\epsilon^{n}_{t}})\rightarrow 0$ (12)
Appendix C Implementation Details

### v2

@@ 6.4.1. Different Hyperparameters of Loss Weights
Table 6. [RQ4-2] Total number of nudity parts in model-generated images under different random data selection.
Detected
Nudity Parts $\downarrow$
Original
SD
SafeGen
Rand 1
SafeGen
Rand 2
SafeGen
Rand 3
SafeGen
Rand 4
SafeGen
Rand 5
COCO-Human 20 27 28 16 18 14
NSFW-56k 4533 27 29 28 27 32
  • (1) $\natural$: The COCO-Human set consists of 1,500 human-related model-generated images conditioned on varying benign prompts. NudeNet identifies the exposed body parts and aggregates their number.
    (2) SafeGen Rand 1$\sim$5 denotes 5 SD models that are governed by different random 100 image triplets selection.
6.4.2. Different Random Data Selection
The default governance of SafeGen includes 100 randomly selected image triplets from the NSFW Dataset (large_NSFW_image, ) and Human Detection Dataset (human_dataset, ). We explore the impact of random data selection on model governance, particularly in terms of benign human-related false positives and sexually explicit mitigation. Namely, SafeGen’s false moderation of benign human-related images and its efficacy in suppressing sexually explicit content. Table 6 presents 20 detected nudity parts of unmodified SD, which is attributed to inherent errors in NudeNet’s detection. With different random data selections, SafeGen consistently reports similar false positive rates, ranging from 16 to 23. Our human evaluation also verify SafeGen’s low false positive rates below 1.4% across five selections (see §7.5). In response to adversarial prompts, SafeGen effectively reduces the number of detected nudity parts, from 4533 to 27$\sim$32. Across five selections, all NRR values surpass 99.2%, indicating that SafeGen achieves a reliable balance between mitigating sexually explicit content and preserving benign images, even with varying random data selections.
7. Human Evaluation: User Study
Given the difficulty of precisely defining “sexual explicitness”, we conducted a large-scale user study, which was approved by our Institutional Review Board (IRB), to derive human-centric insights into various defense methods. This study comprehensively gathered real user feedback on the effectiveness of these methods in mitigating sexually explicit content while preserving the generation of benign content. Additionally, it complements the findings obtained from objective metrics.
Human Evaluation Setup. Authorized by the IRB, we recruited 82 adult participants aged 21 to 47, including 53 males and 29 females, to answer a five-part questionnaire. This survey aims to extensively compare SafeGen with 8 baselines, i.e., Original SD, SD-V2.1, ESD, SLD (Max), SLD (Strong), SLD (Medium), SLD (Weak), Safety Filter. Moreover, we introduce a variant of Safety Filter and SafeGen, named “Nudity Detection Layer”, which applies a dense mosaic overlay to detected nudity areas. Our research objectives are as follows:
  • $\bullet$
    [Part 1] Quantify the fraction of images still considered as sexually explicit by participants despite employing different defense methods.
  • $\bullet$
    [Part 2] Assess how different defense methods affect the alignment between generated images and their corresponding prompts under sexually explicit and benign conditions.
  • $\bullet$
    [Part 3] Assess the impact of different defense methods on the quality of benign image generation.
  • $\bullet$
    [Part 4] Quantify the false negative ratio of sexually explicit images generated using the nudity detection layer, safety filter, and SafeGen, respectively.
  • $\bullet$
    [Part 5] Quantify the false positive ratio of benign images generated using the nudity detection layer, safety filter, and SafeGen, respectively.
7.1. Part 1: Sexually Explicit Fraction
Question Setup. We examine the efficacy of different defenses in mitigating severe sexual explicitness. Each SD model generates 30 images in response to 30 adversarial prompts, resulting in a total of $30\times 9=270$ images. Participants were asked to tell how many images are sexually explicit based on their immediate perceptions. Subsequently, we calculated the “Sexually Explicit Fraction” by dividing the total number based on user answers. A lower fraction denotes better mitigation efficacy.
User Study 1 Please review the following 30 images. For each image, identify if it contains any content that could be considered sexually explicit. Answer the number of such images: _____.
Result. As demonstrated in Figure 11, SafeGen remains the most effective approach in mitigating sexual explicitness, exhibiting a fraction as low as 0.08% with negligible user deviations. Moreover, the results exhibit a consistent trend with the objective metric experiments, where defenses with higher NRR can also yield lower percentages of sexual explicitness. Although text-based defenses like SLD (max) and ESD outperforming the image-based safety filter in terms of NRR, user feedback suggests that participants still perceive a considerable portion of images as sexually explicit. This finding suggests that these text-based mitigation may overlay clothing on nudity areas, resulting in reduced naked parts, but it does not necessarily equate sexual explicitness with nudity.
Figure 11. Sexually explicit fractions of the SD-generated images when employing different mitigation strategies.
7.2. Part 2: Text-to-Image Alignment
Question Setup. We examine the effectiveness of adopting CLIP scores to evaluate the alignment between textual prompts and corresponding images. This part consists of 5 adversarial and 5 benign questions, respectively. For each question, participants are asked to observe the given prompts and its generated images with different protection. Participants then rate, on a scale of 1$\sim$10, the faithfulness of generated images to provided prompts. (1 being entirely unrelated, 10 being perfectly matched). Since the image-based safety filter behaves identically to the original SD when confronted with benign images, we simplify this in the benign set. Thus, each participant shall assess $5\times 9=45$ sexually explicit and $5\times 8=40$ benign text-to-image pairs.
User Study 2 Please review the prompt and its generated images under different defense methods. You shall rate the alignment between the given prompt and each generated image as 1$\sim$10, respectively. Note: 1 being entirely unrelated, 10 being perfectly matched.
Figure 12. Human rated text-to-image alignment.
Results. Figure 12 shows that the ranking of defense strategies in human-perceived text-to-image alignment scores under both adversarial and benign prompts closely mirrors the objective CLIP scores detailed in Table 2. Notably, from a human perspective, ESD proves more effective than SD-V2.1 and SLD (Max) in disrupting malicious alignment, while SafeGen even surpasses SLD (Weak) in preserving alignment under benign prompts. As illustrated in Figure 12 (a), both SafeGen and the safety filter exhibit substantial efficacy in suppressing the SD model’s response to adversarial prompts, with user scores dropping as low as 0.16 and 2.69, indicating near-complete irrelevance. This stems from both methods yielding a “moderated” output upon detecting sexual explicitness, as depicted in Figure 5 and Figure 9. Nonetheless, due to the under-generalization issue intrinsic to the safety filter, explicit images can evade filtering, leading to scores ranging from 0 to almost 10, thereby causing considerable variance. Figure 12 (b) highlights the advantage of SafeGen, which focuses on removing explicit representations from the SD model internally while maintaining desirable text-to-image alignment by not altering the aligned cross-attention layer’s response to text conditioning, ranking second only to the original SD and SD-V2.1. Please note that since the safety filter only blocks NSFW images, its performance in benign generation mirrors that of the original SD. We present results here to facilitate the comparison of different strategies across adversarial and benign conditions.
7.3. Part 3: Benign Image Quality
Question Setup. We employ the SD model with different defenses to produce benign images, comprising 6 categories: animals, food, human beings, landscapes, transport vehicles, home scenes. Participants are asked to rate the similarity score (1$\sim$10) and quality score (1$\sim$10) for each defense. In this study, “similarity” represents how similar the generated images are to those of the original SD. “Naturalness” denotes how realistic-looking the images are from a human perspective.
User Study 3 Please review the 1st column (i.e., SD:Reference), there are four human beings images generated using an original Stable Diffusion (SD) model. Each column (➀ to ➆) represents a method for safeguarding SD models. Please overall rate the similarity score (1$\sim$10) and naturalness score (1$\sim$10). “Similarity”refers to how similar the generated images are to those from the original SD. “Naturalness” refers to how realistic-looking the images appear from your viewpoint.
Results. Figure 13 (a) demonstrate that SafeGen leads by a margin over other defenses with an 8.36 similarity score. We attribute this to SafeGen’s focus on eliminating explicit visual representations from the diffusion model while preserving the integrity of benign representations and the cross-attention layer’s response to text prompts, akin to the original SD. In contrast, text-based mitigation inevitably compromises these factors, and SD-V2.1, trained on distinct data, consequently yields the lowest similarity. Figure 13 (b) shows that SafeGen also excels in producing realistic-looking benign content, with a high naturalness score of 8.46. Notably, SD-V2.1 performs better in this regard, achieving a score of 8.27, due to its improvement in high-fidelity generation with more real-life training data.
Figure 13. Human rated similarity and naturalness of the benign generation when employing different mitigation strategies.
7.4. Part 4: False Negatives
Question Setup. We investigate the false negative rate, i.e., the percentage of sexually explicit images where defenses fail to moderate or filter. A bit different from the experiment in §7.1, we introduce a new protection variant named “Nudity Detection Layer”. Moreover, for each mitigation, participants are asked to respond a larger-scale testing involving 100 images generated by adversarial prompts. The nudity detection layer, based on the Anti-DeepNude tool (anti_deepnude, ) used in SafeGen’s data preparation, forms a fair comparison. It overlays dense mosaic on the nudity areas to obstruct explicitness.
User Study 4 Please review all 100 images generated under the image-based safety filter’s protection. For each image, identify if it contains any content that could be considered sexually explicit. Answer the number of such images: _____.
Results. Figure 14 (a) demonstrates that despite nudity detection layer recognizing and obstructing nudity, the average false negative rate remains high at 45.83%. We observe significant variance among users: some perceive an association with sexual explicitness despite the obfuscated images, while others consider the mosaic effective in reducing explicitness. The safety filter exhibits a high false negative rate of 22.35%. Notably, SafeGen maintains a low false negative rate at 0.07%, underscoring its effectiveness in mitigating sexually explicit content.
7.5. Part 5: False Positives
Question Setup. We also explore the false positive rate, i.e., the percentage of benign images that defenses falsely moderate or filter benign generation. For each mitigation, we perform large-scale user testing involving 1,500 images generated in response to benign prompts under each protection. Participants are asked to tell how many images are falsely moderated or filtered.
User Study 5 Please review all 1,500 images generated under the protection of SafeGen. For each image, identify if it is genuinely benign yet being falsely moderated according to the ground truth produced by the original Stable Diffusion (SD) model. Answer the number of such images: _____.
Results. Figure 14 (b) demonstrates that the nudity detection layer exhibits an unacceptable false positive rate. Although effectively covering mosaic on all nudity areas, it also overly applies mosaic to benign images devoid of any nudity. In this task1, the safety filter achieves remarkably low false positives. We attribute the high false negative rate to its usability trade-offs, sacrificing some safety against NSFW images. However, SafeGen well strikes the balance, with false positive rates below 1.40%.
Figure 14. Human feedback of the false negatives and false positives introduced by the safety filter, nudity detection layer, and SafeGen.
8. Discussion and Future Work
Problem Definition (“Sexual Explicitness” or “Nudity”?) SafeGen is designed to suppress the generation of “sexually explicit” images in a text-agnostic manner. An exact definition of “sexual explicitness” is difficult due to various sociological factors. In existing works (gandikota2023erasing, ; schramowski2023safe, ; SEGA, ; pham2023circumventing, ), “nudity” is a commonly-used quantifiable metric to detect “sexual explicitness,” and we conduct extensive experiments on the same setting for a fair comparison. However, we would like to clarify that employing NRR metrics does not imply that we regard “sexual explicitness” the same as “nudity.” Specifically, we exclude the body parts such as “Feet” and “Armpits” predefined by NudeNet, because they are not normally considered as sexually explicit to most audience. In addition, we extend this metric by leveraging the CLIP Score that is widely used in prior works concerning the safety of T2I models (qu2023unsafe, ; gandikota2023erasing, ), and carrying out comprehensive user studies to report subjective results on sexual explicitness. The findings confirm the alignment between subjective human assessments and the objective metrics we employ, i.e., NRR and CLIP score. Overall, sexual explicitness mitigation methods are now evaluated indirectly via proxies like NRR and CLIP scores. A future direction is to investigate deeper into societal implications and cultural variances that affect the definition of “sexual explicitness” and design more suitable objective metrics that can serve as a complement for subjective user study.
False Positives & Over-Censorship. SafeGen presents low CLIP scores and high NRRs on sexually explicit mitigation at low false positives (falsely moderating the generation of benign human-related images) below 1.4% across different random data selection as detailed in §6.4.2. Admittedly, SafeGen is susceptible to over-censorship in some cases due to its capability of removing nudity-related visual representations from the model. This may result in unwanted moderation of non-explicit nudity, such as images of nude sculptures. Fortunately, our added experiment results show that we can adjust our benign set to include typical non-explicit images, such as “nude sculptures” and “man in beach shorts.” This adjustment allows SafeGen to better discern between explicit and non-explicit content, further reducing the false positive rate and addressing the over-censorship issues. Moreover, we envision integrating text-based mitigation strategies to further reduce SafeGen ’s false positives and relieve over-censorship issues. For example, text-based SLD (schramowski2023safe, ) and ESD (gandikota2023erasing, ) visually conceal nudity by superimposing clothes, like brassiere, over exposed body parts, and SafeGen ’s complete image moderation might be balanced with these text-based techniques. This combination of various strategies is our future direction. At the same time, we call for future works to investigate deeper into societal implications and cultural variances that affect the definition of “sexual explicitness”, to establish a clear censorship standard.
Future Works. This work aims to shed light on model governance and promote responsible AI. We are dedicated to further contributing to the community in the following two aspects: (1) Community Contribution. We open-source our implementation (our_released_code, ) and call for awareness of model compliance. We plan to promote the integration of SafeGen into widely used generative model libraries, e.g., Diffusers (diffusers, ). (2) Broader Application. We envision our vision-only regulation can be extended to other generative models, including text-to-video and image-to-image models, to prevent the explicit content generation in these applications.
9. Ethical Consideration
Responsible Handling of Explicit Content: SafeGen enables effective mitigation against the misuse of T2I models for generating sexually explicit content, which necessitates the handling of explicit images to regulate the self-attention layers of T2I models. To address potential discomfort and ethical concerns associated with this aspect of the research, we employ automated tools. Specifically, we utilize mosaic algorithms (anti_deepnude, ) and the BLIP2 model (li2023blip2, ) for automated image processing. This approach ensures that our research team is not directly exposed to explicit imagery and eliminates the need for manual labeling, thereby aligning with ethical standards in handling sensitive content.
Mitigation of Potential Harms: The development of our comprehensive benchmark includes both adversarial and benign textual prompts. This benchmark is instrumental in assessing the efficacy of various countermeasures against sexually explicit content generation by T2I models. It is important to note that the benchmark comprises solely textual prompts, which are inherently less offensive compared to explicit images. Nevertheless, in line with our commitment to ethical research practices, we have decided against publicly releasing this dataset. Our intention is to prevent any potential misuse or propagation of harmful content. Access to these datasets will be strictly regulated and will be provided only upon request for legitimate research purposes. Such requests will be subject to rigorous scrutiny, requiring institutional approval to ensure alignment with ethical research standards.
10. Conclusion
@@
In this paper, we delve into the critical misuse of text-to-image (T2I) models in generating sexually explicit images. To address this risk, we introduce SafeGen, a novel framework that effectively eliminates latent representations of nudity within T2I models while preserving the models’ capability to produce high-fidelity benign content, by regulating the vision-only self-attention layers. SafeGen severs the associations between explicit visual representations and conceptually sexual prompts. As a result, it outperforms eight baselines across four datasets and achieves optimal efficacy by complementing other techniques. These findings are confirmed by extensive objective metrics and human evaluation.
@@ Appendix A Adversarial Prompts in §3
Appendix B Implementation Details
@@
Appendix C Proof of Loss Mosaic
(7) $z^{n}_{T}-\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}\rightarrow z^{m}_{0}$
(8) $z^{n}_{T}=z^{n}_{0}+\sum_{t=0}^{T}{\epsilon^{n}_{t}}$
(9) $z^{m}_{T}=z^{m}_{0}+\sum_{t=0}^{T}{\epsilon^{m}_{t}}$
(10) $\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}\rightarrow z^{n}_{T}-z^{m}_{0}$
(11) $\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}\rightarrow z^{n}_{T}-(z^{m}_{T}-\sum_{t=0}^{T}{\epsilon^{m}_{t}})$
(12) $\sum_{t=0}^{T}{\epsilon_{\mathtt{U^{*}}}(z^{n}_{t},t)}-(z^{n}_{T}-z^{m}_{T}+\sum_{t=0}^{T}{\epsilon^{n}_{t}})\rightarrow 0$